OpenStudy (anonymous):

solve and show steps 5(1.06^2x+1)=11

OpenStudy (anonymous):

is this \[5\times 1.06^{x+1}=11\]?

OpenStudy (anonymous):

close its 5 X 1.06^2x+1=11

OpenStudy (anonymous):

the +1 is out at the end?

OpenStudy (anonymous):

its 1.06 raised 2x+1

OpenStudy (anonymous):

\[5\times (1.06)^x+1=11\] \[5\times (1.06)^x=10\] \[(1.06)^x=2\] \[x=\frac{\ln(2)}{\ln(1.06)}\]

OpenStudy (anonymous):

thats incorrect

OpenStudy (anonymous):

\[5\times 1.06^{x+1}=11\] \[1.06^{x+1}=\frac{11}{5}=2.2\] \[x+1=\frac{\ln(2.2)}{\ln(1.06)}\] \[x=\frac{\ln(2.2)}{\ln(1.06)}-1\]

OpenStudy (anonymous):

you are forgetting the 2

OpenStudy (anonymous):

5(1.06^2x+1)=11

OpenStudy (anonymous):

the answers should be 6.2657 but i dont know how to get it

OpenStudy (anonymous):

5(1.06^2x+1)=11 =5*1.06^2x+5=11 =5*1.06^2x=11-5=6 =1.06^2x=6/5 Take log of both sides 2xlog(1.06)=log(6/5) 2x=log(6/5)/log(1.06)=3.129 x=1.5645

OpenStudy (anonymous):

i'll ask my teacher in class today

OpenStudy (anonymous):

http://openstudy.com/?F1235834802255UAVKIJ=_#/updates/4ee2419be4b0a50f5c5624fa here i asked and this girl answered it right. i had to do this problem in class.

OpenStudy (anonymous):

so it's right. :)